Solve 2sin3x-1=0
when -180 <= x <= 180
show all work
>>>>.............
If cos x = -0.5 find x when
when -180 <= x <= 180
show all work
>>>> ...........
I asked for this before but I did not get it >> solve this : 2sin3x=1 when
0 < or equal to x < or equal to 90''.
I need the explanation please
and thank you so much

Answer by stanbon(75887)

Solve 2sin3x-1=0
when -180 <= x <= 180
show all work
---
sin(3x) = 1/2
The reference angle is pi/6 and sin is positive in Q1 and QII
-------
3x = pi/6 or 3x = (5/6)pi
x = pi/18 or x = (5/18)pi

>>>>.............
If cos x = -0.5 find x when
when -180 <= x <= 180
show all work
---
The reference angle is pi/3 and cos is negative in QII and QIII
x = (2/3)pi or x = (4/3)pi

>>>> ...........
I asked for this before but I did not get it >> solve this :
2sin3x=1 when
0 <= to x <= 90 degrees
---
sin(3x) = 1/2
reference angle is 30 degrees and sin is positive in QI and QII
3x = 30 degrees
x = 10 degrees
=====================
Cheers,
Stan H.